transform the given differential equation or system into an equivalent system of first order differential equation

x+3x²+48-2y=0 y+24-3x+y = cost
Answer #1

x + 2 x +40-2y =0 9 + 2y - 33 + 5y = cos? Now = 2 -32, -ua+39, - 29, +31. - sg, & cost 0 0 0 0 0 0 U